A 6ft man walks away from a 12ft lamp post at the rate of 4ft/sec. How fast is his shadow lengthening when he is 21ft from the post? 
A. 3ft/sec
B. 4ft/sec
C. 5ft/sec
D. 6ft/sec

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii 
Let the length from pole to man be x 
Length of shadow be y 
From similarity of triangle 
find the distance x 
x = 10.5 
Total distance of shodow from pole = 21 + 10.5 = 31.5 ft
Rate of shadow = distance / time 
for travelling 4 ft – it takes 1 sec.
1 ft – ¼ sec
21 ft – 21/4 
31.5*4/21 = Rate of shadow = 6ft /sec, 
D is correct one.